OpenStudy (anonymous):

Help! :3 Follow and medal if you help: 5t squared + 5s - (3t squared - 6s + 4)

OpenStudy (arindameducationusc):

you there?

OpenStudy (anonymous):

Yes

OpenStudy (arindameducationusc):

5t^2-3t^2=2t^2 5s-6s=-1s

OpenStudy (arindameducationusc):

so equation we get is 2t^2-s-4=0

OpenStudy (anonymous):

Thank you :3